OpenStudy (anonymous):

Could somebody please check if I am doing this right. sketch the graph y=x^4 -4x^2 + 4

OpenStudy (anonymous):

When x=0 , y= 4 so we get point (0,4) Y= x^4 -4x^2+4 becomes x^2(x^2-4) +4 = 0

OpenStudy (anonymous):

So far am I right????

OpenStudy (amoodarya):

hint : y= (x^2-2)^2

OpenStudy (amoodarya):

y= (x-sqrt(2)^2(x+sqrt(2))^2

OpenStudy (amoodarya):